OpenStudy (anonymous):

What is the value for x in the proportion x+8/5x- 2=3/8?

OpenStudy (anonymous):

Solve for x: -2+(13 x)/5 = 3/8 Put the fractions in -2+(13 x)/5 over a common denominator. Put each term in -2+(13 x)/5 over the common denominator 5: -2+(13 x)/5 = (-10)/5+(13 x)/5: (-10)/5+(13 x)/5 = 3/8 Combine (-10)/5+(13 x)/5 into a single fraction. (-10)/5+(13 x)/5 = (-10+13 x)/5: 1/5 (-10+13 x) = 3/8 Multiply both sides by a constant to simplify the equation. Multiply both sides of (13 x-10)/5 = 3/8 by 5: (5 (13 x-10))/(5) = 5×3/8 Express 5×3/8 as a single fraction. 5×3/8 = (5×3)/8: (5 (13 x-10))/5 = (5×3)/8 Cancel common terms in the numerator and denominator of (5 (13 x-10))/5. (5 (13 x-10))/5 = 5/5×(-10+13 x) = -10+13 x: -10+13 x = (5×3)/8 Multiply 5 and 3 together. 5×3 = 15: 13 x-10 = 15/8 Isolate terms with x to the left hand side. Add 10 to both sides: 13 x+(-10+10) = 15/8+10 Look for two terms that sum to zero. -10+10 = 0: 13 x = 15/8+10 Put the fractions in 15/8+10 over a common denominator. Put 15/8+10 over the common denominator 8. 15/8+10 = 15/8+(8×10)/8: 13 x = 15/8+(8×10)/8 Multiply 8 and 10 together. 8×10 = 80: 13 x = 15/8+80/8 Add the fractions over a common denominator to a single fraction. 15/8+80/8 = (15+80)/8: 13 x = (15+80)/8 Evaluate 15+80. 15+80 = 95: 13 x = 95/8 Solve for x. Divide both sides by 13: x = 95/(8×13) Multiply 8 and 13 together. 8×13 = 104: Answer: | | x = 95/104

OpenStudy (anonymous):

@ zkrup bro i think ur ans is nt appropriate..... we will go for cross multiplicatiom i.e 8x+64=15x-6 70=7x x=10

OpenStudy (anonymous):

10 is satisfying this equation